OpenStudy (anonymous):

What is the slope of the graph of 8x – 4y = 11?

OpenStudy (anonymous):

2 write the equation in this format: \[y=mx+c\]

OpenStudy (luigi0210):

Put it in slope intercept form: -4y=-8x+11 divide: y=2x-11/4 so it is 2

OpenStudy (anonymous):

delete wat

OpenStudy (anonymous):

I thight it was -2

OpenStudy (jhannybean):

\[\large 8x-4y=11\] isolate the y by subtracting 8x from both sides. \[\large -8x+8x-4y=11-8x\]simplify\[\large 4y=-8x+11\]divide by 4 toput this equation in slope intercept form. \[\large y=mx+b\]\[\large y=-\frac{8}{4}x+\frac{11}{4}\] the reduce and you'll find the slope using the slope intercept form.

OpenStudy (luigi0210):

lost your negative there

OpenStudy (jhannybean):

xD I did :( y=-8/-4 -11/4 **

OpenStudy (jhannybean):

-_- -8/-4x*

OpenStudy (anonymous):

huh? so 1/2, -2, 2, - 1/2

OpenStudy (luigi0210):

Look at her long and detailed explanation.. :P

OpenStudy (luigi0210):

It is 2 bud

OpenStudy (jhannybean):

I made a booboo. \[\large -4y=-8x+11\]\[\large y=\color{red}{\frac{-8}{-4}}x-\frac{11}{4}\]simplifying you would get your slope. the highlighted part.

OpenStudy (anonymous):

She still did it right bro, but thx